calculate; without balancing following equation , volume of chlorine NTP produced from 50g of sodium chloride.
NaCl+MnO2+H2SO4=NaHSO4+MnSO4+Cl2+H2O

Grade:9

1 Answers

Arun
5 years ago
Chlorine is produced in industries through electrolysis of either Sodium Chloride, brine solution (NaCl) or Potassium chloride solution (KCl)
The reaction for the electrolysis of Brine is as follows:
2 NaCl + 2 H₂O → Cl₂ + H₂ + 2 NaOH
Calculate how many moles 50g NaCl has
moles = mass/molar mass
(molar mass NaCl = 58.5)
moles = 50/58.5
          = 0.855 moles
The mole ratio between NaCl and Cl in the equation above is 2:1
That means moles of chlorine for 50g of NaCl = 0.855/2 = 0.4275 moles
Calculate the volume of the Chlorine gas using its moles:
Ideal gas law states that 1 mole of an ideal gas occupies 22.4 liters by volume.
If 1 mole = 22.4 liters
Then 0.4275 moles = 22.4 × 0.4275/1
                               = 9.573 liters of chlorine.